Hi, my name is Kiara and I'm having some problems with my homework I will aprreciate all you're help .
solve the given formula for the variable shown in color. State the restrictions, if any, for the formula obtained to be meaningful.
p=2(l + w) : w
Can you explain to me how to solve it??
Thank you, and have a great day..

Answer by stanbon(75887)

p=2(l + w) : w
Can you explain to me how to solve it??
--------------------------
Divide both sides of the equation by 2 to get:
p/2 = l+w
---------------
Subtract "l" from both sides of the equation to get:
w = (p/2)-l
=======================
Cheers,
Stan H.